OpenStudy (anonymous):

Bungles the Clown is giving away balloons to the children at the school carnival. He has 10 red balloons, 7 green balloons, 11 blue balloons, 14 yellow balloons, and 8 purple balloons. Bungles randomly chooses a balloon to give to each child. What is the probability that Bungles will select a yellow balloon?

OpenStudy (anonymous):

number of yellow balloons divided by the total number of balloons

OpenStudy (anonymous):

14/total possibilites =14/(10+7+11+14+8) =14/50 = 7/25